OpenStudy (anonymous):

solve: log10^z+log10^(z+3)=1?

OpenStudy (anonymous):

log10^z+log10^(z+3) =1 =>zlog10 +(z+3)log10 = 1 =>z +z+3 =1 can you find z now??

OpenStudy (anonymous):

Yes, Thanx!

OpenStudy (anonymous):

welcome